The solubility product of FeCO3 is given by

Ksp = 2.1 x 10-11 M2

What is the maximum solubility of FeCO3 in pure water?

S = ___ M
Answer #1

At equilibrium:

FeCO3 <----> Fe2+ + CO32-

   s s

Ksp = [Fe2+][CO32-]

2.1*10^-11=(s)*(s)

2.1*10^-11= 1(s)^2

s = 4.583*10^-6 M

Answer: 4.6*10^-6 M